An investor invested a total of $2,600 in two mutual funds. One fund earned a 4% profit while the other earned a 2% profit. If the investor's total profit was $84, how much was invested in each fund?


Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 2.00% per annum ------------ Amount invested = y
2600
Interest----- 84.00

Part I 4.00% per annum ---x
Part II 2.00% per annum ---y
Total investment
x + 1 y= 2600 -------------1
Interest on both investments
4.00% x + 2.00% y= 84
Multiply by 100
4 x + 2 y= 8400.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -10400.00
Add this to (2)
0 x -2 y= -2000
divide by -2
y = 1000
Part I 4.00% $ 1600
Part II 2.00% $ 1000

CHECK
1600 --------- 4.00% ------- 64.00
1000 ------------- 2.00% ------- 20.00
Total -------------------- 84.00